OpenStudy (anonymous):

What are the solutions for x in the proportion 4 times x minus 3 all over x equals 2 times x all over 2.

OpenStudy (anonymous):

\[\frac{ 4x-3 }{ x } = \frac{ 2x }{ 2 }\]

OpenStudy (anonymous):

Cross multiply? 2x^2 = 8x - 6 ?

OpenStudy (anonymous):

@So1D

OpenStudy (anonymous):

Lol Be Patient With Me Im Slow With Math :D

OpenStudy (anonymous):

That's fine ;b

OpenStudy (anonymous):

@FutureMathProfessor Maybe you can help?o;

OpenStudy (anonymous):